Systems Staff Meeting Notes
November 9, 2007
November 9, 2007
Present: Thomas Bielicke, Cynthia Bowker, Will Chaney, Logan Clapp, Debra Fetzer, Winston Harris, Bobbie Meng, Tom Minton, Chris Nicolich, Amy Polk, Mark Sullivan, Stephen WilliamsAnnouncements
Review Minutes of 10/26/07
- It's time for vacation leave sell-back for TEAMS employees with enough hours. December leave cash-out: 16 hours.
- Dean Russell will be returning to the Systems staff meeting on 1/18/08.
- Second half of the pay raise will start in the 11/21/07 paycheck. $1000 bonus in today's paycheck.
Ongoing Systems Updates
- Remove sentence in Software #2.
- Web server had a drive failure last night. Reboot caused problems. We're in the process of rebuilding a new drive. Once the SAN is up, we can move the Web server.
- Celebration was sparsely attended. We received a reward from CNS - it is hanging up in room 10.
Software Projects Status
- Smathers NT2R1 - problem this morning - failed over to NT2R. It has now been restarted and seems okay.
- Campus network outage caused by a Ricoh copier flooding the network.
- We had a Xerox copier that needed many updates - we were advised by CNS. It had a hard drive failure - had to rebuild and bring it back up. They need to be updated before they are brought back on the network.
- Education server - 2K3Edu failed, it is being rebuilt.
- Master browser issues - Will is tinkering with it.
- McAfee updates - Will has been working on it - it is down to problems on individual machines. They need to install the framework from the server. New version of EPolicy is out; need to test then push.
- SAN is installed, all drives are configured. We are currently discussing the best way to divide up the storage space. ESX, VM host server software, is installed. Logan is working on it.
- SMS - Amy is educating herself on it.
- Will removed 4.53G of MP3 files from NT2R. SMS will scan for MP3s.
- 3480 cartridges are being converted to the virtual server at CNS. Debra met with Scott Crumpton and he modified the menu program in TSO to read and map the tapes. Debra verified the program will work.
- Central Exchange server - they will start converting those already in line on the MLK weekend. Retention policy - huge change for users - will purge your stuff automatically. You will be able to have a managed folder for archival purposes. Deleted items will be deleted in 7 days. Theoretically, mailbox size will be unlimited. We'll have to inform users. We will be one of three tiers of support. There will be plenty of training opportunities. Web access will be the same. They will also provide Blackberry and Palm support. Log-in disclaimer - there is no good way to track users. We may want to create a program to track. Logan is working with John Sawyer using Syslog for all clients and servers.
- Backup Exec upgrade - Cynthia completed on Thursday. We are on the new version. It will be on continuous protection mode for Exchange mailboxes.
- Lenovo tech support - put Darren's number in a folder on the server. Discussion about putting contact numbers on a secure Web page. Good idea.
- Blog server is down. Logan is moving all VMs to a dedicated server. It won't be back up until next week. Tom asks that the VM Web server be put back up.
- Grover - ongoing. Working on tickets.
- Tracking database - new version of the importer.
- LibQual survey is finished for the year.
- Brian and Trudi have asked for additional ITS reports.
- Greenstone - optimizing UFDC front end. Small features added. It has been up and down a bit over the last few days. Need to turn the loader back on. Need to change passwords.
- Barcode prefix - per Lori Driscoll, it is not a high priority. She will advise as it comes up.
- Winston started up a project to clean-up PeopleSoft - Library patrons with no affiliations - creating a way to track. Old accounts have not been deleted. Will devise a way to get those accounts to expire, with renewal possibilities.
Systems Security Management
- Tom had a request from Laurie Taylor (DLC) to use server-side includes. She is willing to instruct the rest of the Web people on how to use it.
- Complaints that Google is finding incorrect Web sites for the Libraries. This is mainly due to out-of-date Web pages. Someone is asking for a 301 redirect on the server to the new Career Development manual. They really need to just remove the old sites. Will to discuss the issue.
- Meeting held.